231 milliwatts in scientific notation without prefixes. (a) 9.231 x 10-3 Watts (b) 9.231 x 10-4 Watts (c) 9.231 x 107 Watts (d) 9.231 x 10-7 Watts (e) 9.231 x 101 Watts 6. Rewrite 53657 nanoseconds in scientific notation without prefixes. (a) 5.3657 x 1013 seconds (b) 5.3657 x 104 seconds (e) 5.3657 x 10- seconds() 5.3657 x 10-13 (d) 5.3657 x 10-9 seconds 7. Rewrite 5.32 milligrams in scientific notation without prefixes. (a) 5.32 x 10-5 grams (b) 5.32 x 103 grams (c) 5.32 x 10-1 grams (d) 5.32 x 10° grams (e) 5.32 x 103 grams 8. Rewrite 88900 megahertz in scientific notation without prefixes.
Answer #1

5)D) 0.0009231 milliwatts = 9.231 x 10-4 milliwatts = 9.231 x 10-7 watts

6)C) 53657 nano seconds = 5.3657 x 10+4 nanoseconds = 5.3657 x 10-5 seconds

7)C) 5.32 milligrams=5.32 x 10-3 kilograms

8) 88900 megahertz=8.89 x 10+4 megahertz= 8.89 x 10+10 hertz